# Enhancement of temperature of quantum anomalous Hall effect in two-dimensional germanene/magnetic semiconductor heterostructures

### v1

@@ I Introduction
Quantum anomalous Hall effect (QAHE) is one of the most exciting phenomenon of matter and is characterized by an integer Chern number $\mathcal{C}$ [1, 2]. A key manifestation of QAHE is the dissipationless chiral edge current. It has the potential to revolutionize future low-power quantum electronic devices. Therefore, much theoretical and experimental research has been conducted to achieve QAHE at high temperatures [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Various methods are being explored to achieve QAHE. Introducing ferromagnetic order in topological insulators (TIs) is a primary method for achieving QAHE. The first observation of QAHE occurred in magnetically doped TIs, i.e., Cr-doped (Bi,Sb)2Te3 thin films at 30 mK [3]. Subsequent research has focused on increasing the QAHE temperature. For example, the anomalous Hall resistance $\rho_{xy}$ reaches a maximum value of $0.95h/e^{2}$ in a 10 quintuple-layer (QL) thin film of Cr0.1(Bi0.5,Sb0.5)1.9Te3 at 280 mK [8], and a six QL (Cr0.12Bi0.26Sb0.62)2Te3 film exhibits a quantized anomalous Hall resistance $R_{yx}$ up to 0.3 K [10]. To date, the highest reported record QAHE temperature in these doped magnetic TIs is approximately 2 K [11]. Because disordered magnetic impurities can inevitably degrade sample quality, thus limiting the QAHE temperature. Many researchers are investigating intrinsic QAHE materials. In recent years, experimental breakthroughs in QAHE have been achieved in the van der Waals (vdW) layered material MnBi2Te4 [12, 13, 14]. MnBi2Te4 is a layered magnetic TI consisting of Te-Bi-Te-Mn-Te-Bi-Te septuple layers (SLs) with antiferromagnetic ground state characterized by antiferromagnetic interlayer coupling and ferromagnetic intralayer coupling. In five SL flakes, the quantization temperature can be increased to 6.5 K by aligning all layers ferromagnetically with an external magnetic field [12].
@@ III QAHE in germanene/ML Cr2Ge2Te6 heterostructure
The electronic band structures and density of states (DOS) of germanene/ML Cr2Ge2Te6 heterostructure are shown in Fig. 2. With SOC included, the germanene/ML Cr2Ge2Te6 heterostructure exhibits a finite band gap of 46 meV. This indicates that even at room temperature, thermal fluctuations are insufficient to fully populate the band gap. The Dirac cone around the $\Gamma$ point is predominantly contributed by $p_{z}$ orbits of germanene. Notably, this band gap is larger than that of pristine germanene. Since SOC depends on the atomic number $(\sim Z^{4})$ and SOC strength of Te is larger than that of germanene, the observed band gap enhancement is attributed to interlayer coupling and hybridization. To investigate the topological properties, maximally localized Wannier functions (MLWFs) implemented in the Wannier90 code were used to fit our density functional theory (DFT) band calculations. The calculated Chern number for the valence band near the Fermi level is $C=1$, confirming the topologically nontrivial nature of the band structure. Fig. 3(a) shows the calculated anomalous Hall conductance as a function of chemical potential, revealing a quantized Hall plateau at $\sigma_{xy}=\frac{e^{2}}{h}$. The plateau is consistent with the presence of one topologically nontrivial edge state connecting the bulk states in Fig. 3(b).
@@
Due to the crystal field, the $d$ electron configuration of Cr is split into three groups: $a_{1}$ ($d_{z^{2}}$ orbital), $e_{1}$ ($d_{xy}$ and $d_{x^{2}-y^{2}}$ orbitals) and $e_{2}$ ($d_{xz}$ and $d_{yz}$ orbitals). Wannier90 calculations indicate that the main contribution comes from $e_{2}$. The values of $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ for the Cr1-Te1-Cr2 path in ML Cr2Ge2Te6 without and with tensile strain are listed in Table. 1. The resulting superexchange interaction $J_{super}$ for relaxed and strained ML Cr2Ge2Te6 are $0.372/A$ and $0.878/A$, respectively. The ratio of $J_{super}$ for strained to relaxed ML Cr2Ge2Te6 is 2.36, which is comparable to the ratio (1.25) of the DFT-calculated nearest neighbor exchange coupling $J_{1}$ for the strained and relaxed germanene/ML Cr2Ge2Te6 heterostructure. Comparison of the strained and unstrained ML Cr2Ge2Te6 results reveals that the enhancement of the superexchange interaction $J_{super}$ is primarily due to the decreased energy difference $E_{p^{\prime}}-E_{d_{2}}$. This decrease in $E_{p^{\prime}}-E_{d_{2}}$ contributes to the enhancement of both $J_{super}$ and $T_{C}$. The Cr-Te-Cr bond angles are 89.8 and 91.2 for relaxed and strained ML Cr2Ge2Te6, respectively, indicating no significant difference in the hopping parameter $|V_{pd_{1}}|^{2}$ as shown in Table. 1.
@@
Table 1: $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ between $5p$ orbitals of Te and $d_{xz}$, $d_{yz}$ orbitals of Cr for relaxed and strained ML Cr2Ge2Te6 (in units of eV).
Strain = 0%
$p_{z}-d_{xz}$ $p_{x}-d_{xz}$ $p_{y}-d_{xz}$ $p_{z}-d_{yz}$ $p_{x}-d_{yz}$ $p_{y}-d_{yz}$
$|V_{pd_{1}}|^{2}$ 0.071 0.012 0.153 0.100 0.222 0.127
$E_{p^{\prime}}-E_{d_{2}}$ 0.621 0.742 0.815 0.628 0.749 0.822
$J_{super}A$ 0.372
Strain = 2%
$p_{z}-d_{xz}$ $p_{x}-d_{xz}$ $p_{y}-d_{xz}$ $p_{z}-d_{yz}$ $p_{x}-d_{yz}$ $p_{y}-d_{yz}$
$|V_{pd_{1}}|^{2}$ 0.057 0.011 0.122 0.086 0.182 0.127
$E_{p^{\prime}}-E_{d_{2}}$ 0.729 0.670 0.743 0.730 0.670 0.743
$J_{super}A$ 0.878
@@ IV QAHE in germanene/BL Cr2Ge2Te6 and germanene/ML Cr2Si2Te6 heterostructures
V Conclusion

### v2

@@ I Introduction
Quantum anomalous Hall effect (QAHE) is one of the most exciting phenomenon of matter and is characterized by an integer Chern number $\mathcal{C}$ [1, 2]. A key manifestation of QAHE is the dissipationless chiral edge current. It has the potential to revolutionize future low-power quantum electronic devices. Therefore, much theoretical and experimental research has been conducted to achieve QAHE at high temperatures [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 21, 22, 23, 24]. Various methods are being explored to achieve QAHE. Introducing ferromagnetic order in topological insulators (TIs) is a primary method for achieving QAHE. The first observation of QAHE occurred in magnetically doped TIs, i.e., Cr-doped (Bi,Sb)2Te3 thin films at 30 mK [3]. Subsequent research has focused on increasing the QAHE temperature. For example, the anomalous Hall resistance $\rho_{xy}$ reaches a maximum value of $0.95h/e^{2}$ in a 10 quintuple-layer (QL) thin film of Cr0.1(Bi0.5,Sb0.5)1.9Te3 at 280 mK [8], and a six QL (Cr0.12Bi0.26Sb0.62)2Te3 film exhibits a quantized anomalous Hall resistance $R_{yx}$ up to 0.3 K [10]. To date, the highest reported record QAHE temperature in these doped magnetic TIs is approximately 2 K [11]. Because disordered magnetic impurities can inevitably degrade sample quality, thus limiting the QAHE temperature. Many researchers are investigating intrinsic QAHE materials. In recent years, experimental breakthroughs in QAHE have been achieved in the van der Waals (vdW) layered material MnBi2Te4 [12, 13, 14]. MnBi2Te4 is a layered magnetic TI consisting of Te-Bi-Te-Mn-Te-Bi-Te septuple layers (SLs) with antiferromagnetic ground state characterized by antiferromagnetic interlayer coupling and ferromagnetic intralayer coupling. In five SL flakes, the quantization temperature can be increased to 6.5 K by aligning all layers ferromagnetically with an external magnetic field [12]. Recent studies have revealed the QAHE in moiré materials can be derived from both graphene [21, 22, 23] and transition-metal dichalcogenides (TMDs) [24]. In these systems, QAHE originates from spontaneous valley polarization. However, the typically small band gap of such systems imposes significant limitations on the realization temperature of the QAHE [25, 26].
@@ III QAHE in germanene/ML Cr2Ge2Te6 heterostructure
The electronic band structures and density of states (DOS) of germanene/ML Cr2Ge2Te6 heterostructure are shown in Fig. 2. With SOC included, the germanene/ML Cr2Ge2Te6 heterostructure exhibits a finite band gap of 46 meV. This indicates that even at room temperature, thermal fluctuations are insufficient to fully populate the band gap. The Dirac cone around the $\Gamma$ point is predominantly contributed by $p_{z}$ orbits of germanene. Notably, this band gap is larger than that of pristine germanene. Since SOC depends on the atomic number $(\sim Z^{4})$ and SOC strength of Te is larger than that of germanene, the observed band gap enhancement is attributed to interlayer coupling and hybridization. To investigate the topological properties, maximally localized Wannier functions (MLWFs) implemented in the Wannier90 code were used to fit our density functional theory (DFT) band calculations. The anomalous Hall conductivity is calculated using $\sigma_{xy}=\mathcal{C}\frac{e^{2}}{h}$, where $\mathcal{C}$ is the Chern number defined as
[…]
and ${\bf{\Omega}}(k)$ is the Berry curvature that can be obtained as
${\bf{\Omega}}({\bf{k}})=-\sum_{n<E_{F}}\sum_{m\neq n}2Im\frac{<\psi_{n{\bf{k}}}|v_{x}|\psi_{m{\bf{k}}}><\psi_{m{\bf{k}}}|v_{y}|\psi_{n{\bf{k}}}>}{(\epsilon_{m{\bf{k}}}-\epsilon_{n{\bf{k}}})^{2}},$ (2)
where $\psi_{m(n){\bf{k}}}$ are the Bloch wave functions. $\epsilon_{m(n){\bf{k}}}$ are the eigenvalues and $v_{x(y)}$ are the velocity operators. Such calculation has been implemented in the WannierTools code [45]. The calculated Chern number for the valence band near the Fermi level is $\mathcal{C}=1$, confirming the topologically nontrivial nature of the band structure. Fig. 3(a) shows the calculated anomalous Hall conductance as a function of chemical potential, revealing a quantized Hall plateau at $\sigma_{xy}=\frac{e^{2}}{h}$. The plateau is consistent with the presence of one topologically nontrivial edge state connecting the bulk states in Fig. 3(b).
@@
Table 1: $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ between $5p$ orbitals of Te and $d_{z^{2}}$, $d_{x^{2}-y^{2}}$ orbitals of Cr for relaxed and strained ML Cr2Ge2Te6 (in units of eV).
Strain = 0%
$p_{z}-d_{z^{2}}$ $p_{x}-d_{z^{2}}$ $p_{y}-d_{z^{2}}$ $p_{z}-d_{x^{2}-y^{2}}$ $p_{x}-d_{x^{2}-y^{2}}$ $p_{y}-d_{x^{2}-y^{2}}$
$|V_{pd_{1}}|^{2}$ 1.539 0.439 0.051 8.5E-5 0.026 0.005
$E_{p^{\prime}}-E_{d_{2}}$ 0.430 0.551 0.415 0.317 0.438 0.302
$J_{super}A$ 16.33
Strain = 2%
$p_{z}-d_{z^{2}}$ $p_{x}-d_{z^{2}}$ $p_{y}-d_{z^{2}}$ $p_{z}-d_{x^{2}-y^{2}}$ $p_{x}-d_{x^{2}-y^{2}}$ $p_{y}-d_{x^{2}-y^{2}}$
$|V_{pd_{1}}|^{2}$ 1.407 0.425 0.065 1.5E-4 0.023 0.005
$E_{p^{\prime}}-E_{d_{2}}$ 0.430 0.480 0.516 0.285 0.335 0.371
$J_{super}A$ 18.43
Due to the crystal field, the $d$ electron configuration of Cr is split into two groups: $e_{g}$ ($d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals) and $t_{2g}$ ($d_{xz}$, $d_{yz}$ and $d_{xy}$ orbitals). Wannier90 calculations indicate that the main contribution comes from $e_{g}$. The values of $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ for the Cr1-Te1-Cr2 path in ML Cr2Ge2Te6 without and with tensile strain are listed in Table. 1. The resulting superexchange interaction $J_{super}$ for relaxed and strained ML Cr2Ge2Te6 are $16.33/A$ and $18.43/A$, respectively. The ratio of $J_{super}$ for strained to relaxed ML Cr2Ge2Te6 is 1.13, which is comparable to the ratio (1.25) of the DFT-calculated nearest neighbor exchange coupling $J_{1}$ for the strained and relaxed germanene/ML Cr2Ge2Te6 heterostructure. Comparison of the strained and unstrained ML Cr2Ge2Te6 results reveals that the enhancement of the superexchange interaction $J_{super}$ is primarily due to the decreased energy difference $E_{p^{\prime}}-E_{d_{2}}$. This decrease in $E_{p^{\prime}}-E_{d_{2}}$ contributes to the enhancement of both $J_{super}$ and $T_{C}$. The Cr-Te-Cr bond angles are 89.8 and 91.2 for relaxed and strained ML Cr2Ge2Te6, respectively, indicating no significant difference in the hopping parameter $|V_{pd_{1}}|^{2}$ as shown in Table. 1.
The application of strain could enhance the QAH transition temperature in the germanene/ML Cr2Ge2Te6 heterostructure. Because $T_{C}$ in 2D magnetic semiconductor Cr2Ge2Te6 can be dramatically increased by strain, which has been demonstrated in recent experiment [33]. The fundamental mechanism lies in strain-induced reduction of the energy difference between anions and cations, which contributes to the enhancement of $T_{C}$ by the superexchange mechanism. However, the application of strain simultaneously reduces the band gap of the heterostructure germanene/ML Cr2Ge2Te6 by lowering the conduction bands. Consequently, the upper limit of the QAH temperature in germanene/ML Cr2Ge2Te6 heterostructure is determined by the intersection of $T_{C}$ and band gap curves.
@@ IV QAHE in germanene/BL Cr2Ge2Te6 and germanene/ML Cr2Si2Te6 heterostructures
V Discussion and Conclusion
The QAHE temperature of materials cannot exceed its $T_{C}$ and the topological band gap. For the MnBi2Te4 with band gap of 100-200 meV ($\approx$ 1160-2320 K) and $T_{C}$ of 20 K, the upper limit of the QAHE temperature cannot exceed 20 K. For the germanene/ML Cr2Ge2Te6 heterostructure with biaxial tensile strain, the upper limit of the QAHE temperature cannot exceed 62 K, as shown in Fig. 5. Thus, the higher upper limit of QAHE temperature in germanene/ML Cr2Ge2Te6 heterostructure is expected larger than that in MnBi2Te4.
The realization of the room-temperature QAH state requires two fundamental criteria: (i) a $T_{C}$ exceeding 300 K, and (ii) a topological band gap larger than 300 K ($\approx$ 26 meV). However, in practical materials implementations, these ideal conditions often remain unattainable. A representative example is MnBi2Te4, where the QAH effect was experimentally observed at cryogenic temperatures ($\sim$ 1-2 K), significantly below both its intrinsic $T_{C}$ ($\sim$ 20 K) and the estimated band gap energy (100-200 meV) [12, 13, 14]. The experimental observation deviates from theoretical predictions and can be attributed to two principal factors: (1) extrinsic material imperfections such as structural disorders in samples, and (2) intrinsic magnetic fluctuations occurring below $T_{C}$ that effectively suppress the QAH transition temperature.
The disorder effect can be categorized into two distinct types: magnetic disorder and nonmagnetic disorder. Many studies have revealed that weak magnetic disorder stabilizes the edge states and the QAH conductance, whereas strong disorder drives the system into an Anderson insulating phase [52, 53, 54, 55, 56, 57, 58]. The nonmagnetic disorder induces layer-dependent topological phase transitions. Weak disorder generally enhances the robustness of topological edge states and preserves the QAH conductance. By increasing disorder, the quantization will also be disrupted [59, 60].
We applied in-plane biaxial tensile strain that preserves the space group symmetry of the system. The time-reversal symmetry is broken by the magnetic semiconductor in the heterostructure to realize the QAH state. Strain can effectively enhance Curie temperature $T_{C}$ of some 2D ferromagnetic semiconductors, including Cr2Ge2Se6 [42], MnSe2 [61] and Fe2Cl3I3 [62]. When these materials form heterostructures with topological insulators, it is highly expected that the QAH transition temperatures could be synergistically enhanced due to the strain-enhanced $T_{C}$.
We note that the quantum spin Hall (QSH) effect at room temperature has been demonstrated in Bi4Br4 [63]. In our paper, the QAH effect has been studied, which only has been observed in experiments at very low temperature so far. According to the results in our paper, it will be very interesting to explore the possible high temperature QAH effect in the heterostructures Bi4Br4/FS (FS = ferromagnetic semiconductors).